# RIFT: Repurposing Negative Samples via Reward-Informed Fine-Tuning

### v1

@@ 2 Related Works
Supervised Fine-Tuning (SFT) has become the standard post-training paradigm for adapting pretrained models to specific tasks using high-quality, labeled datasets (Zhang et al., 2023a; Chung et al., 2024a). Although the availability of high-quality instruction-following datasets (Cobbe et al., 2021; Mishra et al., 2022; Zhou et al., 2023; Taori et al., 2023) has significantly enhanced the efficacy of SFT, studies (Dodge et al., 2020; Howard and Ruder, 2018; Ouyang et al., 2022b) indicate that SFT often suffers from overfitting and suboptimal generalization.
To mitigate the challenges of SFT data curation, Reinforcement Learning (RL) has emerged as a powerful alternative for post-training. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022b) and Reinforcement Learning from Verifiable Reward (RLVR) (Guo et al., 2025; Shao et al., 2024) are widely used to align models with human preferences and enhance reasoning, supported by algorithms like DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), GRPO (Shao et al., 2024), and DAPO Yu et al. (2025). However, in contrast to SFT, RL-based training is often more complex, necessitating intricate engineering frameworks (Zheng et al., 2025).
@@ 4.1 Experiment Details
Qwen-2.5-Math-7B Base 54.8 50.3 12.2 16.4 12.1 36.9 20.5 29.0
@@ 5 Conclusion
In this work, we propose RIFT, a simple yet effective post-training framework that leverages the full distribution of self-generated samples. Unlike RFT, which discards valuable negative samples via hard thresholding, RIFT leverages both high- and low-reward trajectories. To ensure optimization stability, we introduce a principled loss formulation that effectively prevents training collapse during reward integration. Extensive evaluation across seven mathematical reasoning benchmarks shows that RIFT consistently outperforms established baselines. This demonstrates that explicitly learning from mixed-quality data, rather than filtering it, allows models to better internalize the structure of correct reasoning and common failure modes. As a robust and data-efficient alignment method, RIFT enables scalable self-improvement without reliance on extensive expert-labeled data.
@@
Appendix A Appendix
@@ A.1 Ablation Study on the Number of Self-Generated Responses
The performance trend is intriguing because the underlying data statistics, reported in Tables 9 and 10, show little variation across $K$: the proportion of positive responses remains stable at around 66.4%, and the fraction of problems with mixed reward signals plateaus near 85% for $K\geq 4$. In other words, simply generating more responses does not significantly alter the overall composition of the training data.
@@ A.2 Exploration Study: RIFT Drives Policy Convergence
As shown in Table 11, the results support the hypothesis that RIFT itself drives capability gains. Across all previously trained models, applying a second round of RIFT yields consistent improvements in Mean@3, outperforming their single-round counterparts: notable gains of +11.5 (SFT), +5.6 (DFT), and +1.2 (RIFT), even when starting from a strong RIFT-initialized policy. Gains diminish as the base policy strengthens, indicating convergence toward a shared high-reward policy. However, the re-application of RIFT primarily boosts Mean@3 rather than Pass@3, which shows little to no further improvement compared to single-round trained RIFT. Together, these findings demonstrate that RIFT acts not as a plug-and-play refiner whose efficacy depends on initialization, but as a self-convergent alignment phase that capable of steering diverse initial policies toward comparable final performance. Tables 12 and 13 provide a detailed statistical overview of the self-generated responses produced by the previously trained SFT, DFT, and RIFT models.
A.3 Proofs and Theoretical Analysis in Section 3
@@
Theorem A.1 (Formal Statement of Theorem 3.4).

### v2

@@ 2 Related Works
Supervised Fine-Tuning (SFT) has become the standard post-training paradigm for adapting pretrained models to specific tasks using high-quality, labeled datasets (Zhang et al., 2023a; Chung et al., 2024a). Although the availability of high-quality instruction-following datasets (Cobbe et al., 2021; Mishra et al., 2022; Zhou et al., 2023a; Taori et al., 2023) has significantly enhanced the efficacy of SFT, studies (Dodge et al., 2020; Howard and Ruder, 2018; Ouyang et al., 2022b) indicate that SFT often suffers from overfitting and suboptimal generalization.
To mitigate the challenges of SFT data curation, Reinforcement Learning (RL) has emerged as a powerful alternative for post-training. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022b) and Reinforcement Learning from Verifiable Reward (RLVR) (Guo et al., 2025; Shao et al., 2024) are widely used to align models with human preferences and enhance reasoning, supported by algorithms like DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024a), GRPO (Shao et al., 2024), and DAPO Yu et al. (2025). However, in contrast to SFT, RL-based training is often more complex, necessitating intricate engineering frameworks (Zheng et al., 2025).
@@ 4.1 Experiment Details
Qwen-2.5-Math-7B Base 92.1 83.6 36.4 42.5 30.0 70.0 48.2 50.7
@@ 4.3 Extending to Reasoner Model
4.4 Generalization to General Preference Alignment
While RIFT is primarily motivated by tasks with verifiable rewards (e.g., mathematical reasoning), we further evaluate its generalizability on broader preference alignment. We fine-tune the Llama3.2-3BGrattafiori et al. (2024) model on the UltraFeedback Cui et al. (2024) dataset and compare RIFT against the standard DPO baseline across three diverse benchmarks: IFEvalZhou et al. (2023b), TruthfulQALin et al. (2022), and HellaSwagZellers et al. (2019).
As shown in Table 8, RIFT consistently outperforms DPO across all three benchmarks, demonstrating its effectiveness beyond verifiable reward tasks. Notably, RIFT achieves the improvement with approximately 50% less GPU memory usage than DPO. We observe that the performance gains on RLHF tasks are more modest compared to mathematical reasoning. We attribute this to the inherent noise in preference datasets (e.g., annotator bias, model-generated preferences), which provide less clean supervision signals than the deterministic rewards available in math tasks. Nevertheless, RIFT’s consistent improvements confirm its robustness even under noisy feedback conditions.
Llama-3.2-3B IFEval TruthfulQA Hellaswag Avg.
Base 12.23 19.42 52.32 27.99
DPO 18.34 22.12 53.74 31.40
RIFT 18.63 24.94 55.89 33.15
Table 8: Zero-shot evaluation on general preference alignment tasks. RIFT outperforms DPO while using $\sim$50% less memory.
5 In-depth Analysis
@@ 5.1 Reward Strategy and Robustness Analysis
Model Method GSM8K MATH Minerva Olympiad AIME24 AMC23 College Avg.
[…]
Qwen-2.5-Math-1.5B (Mean@8) Base 42.6 35.6 9.7 22.6 7.1 31.9 8.2 22.5
SFT 57.0 42.9 9.3 16.1 3.3 21.9 19.9 24.3
DFT 76.8 53.9 15.6 19.1 4.2 25.6 36.4 33.1
RFT 48.8 37.2 13.5 22.5 8.8 33.4 15.2 25.6
DPO 61.8 50.3 11.3 26.7 7.1 41.2 18.6 31.0
SimPO 65.8 50.0 18.0 18.4 5.8 20.6 35.0 30.5
KTO 69.9 58.5 17.2 28.3 8.8 40.9 26.9 35.8
[…]
Qwen-2.5-Math-1.5B (Pass@8) Base 88.0 75.1 32.0 46.5 23.3 67.5 30.1 51.8
SFT 93.6 80.5 29.0 43.7 16.7 60.0 49.0 53.2
DFT 94.5 76.7 38.2 43.0 20.0 55.0 52.4 54.3
RFT 87.0 67.3 30.1 27.5 5.0 39.1 41.1 42.4
DPO 92.9 83.2 33.8 48.7 33.3 72.5 45.4 58.5
SimPO 94.5 81.8 40.8 45.6 16.7 62.5 55.5 56.8
KTO 93.5 82.1 44.9 50.2 30.0 77.5 42.8 60.1
[…]
Table 9: Comparison with SimPO and KTO on Qwen2.5-Math-1.5B. Results show Mean@8 and Pass@8 accuracy (%) across 7 mathematical benchmarks.
Qwen-2.5-Math-1.5B SFT DFT RFT DPO SimPO KTO RIFT
Peak Memory usage (GB) $\downarrow$ 17.95 26.90 18.12 43.31 21.40 42.93 20.10
Average Accuracy (%) $\uparrow$ 24.30 33.10 25.60 31.00 30.50 35.80 37.00
Table 10: Efficiency and performance trade-off against SimPO and KTO on Qwen2.5-Math-1.5B. Results show peak GPU memory usage and average mathematical reasoning accuracy.
@@ 5.2 Computational Efficiency
5.3 Comparison with Efficient Alignment Baselines
To further evaluate the efficiency and effectiveness of RIFT, we extend our comparison to include recent reference-free and low-memory alignment methods, specifically SimPO Meng et al. (2024b) and KTO Ethayarajh et al. (2024). We conduct experiments using the Qwen2.5-Math-1.5B model to assess both performance gains and computational overhead. As illustrated in Table 9 and 10, RIFT achieves a superior trade-off between mathematical reasoning performance and computational overhead. Although SimPO achieves memory savings similar to RIFT by removing the reference model, its performance gains over DPO are marginal, trailing RIFT by 6.5 points. Similarly, while KTO shows a notable improvement over DPO (+4.8 points), it still lags behind RIFT by 1.2 points. Crucially, unlike KTO which maintains a high memory usage, RIFT delivers superior performance with significantly higher resource efficiency.
@@ 6 Conclusion
We propose RIFT, a simple yet effective post-training framework that leverages the full distribution of self-generated samples. Unlike RFT, which discards valuable negative samples via hard thresholding, RIFT leverages both high- and low-reward trajectories. To ensure optimization stability, we introduce a principled loss formulation that effectively prevents training collapse during reward integration. Extensive evaluation across seven mathematical reasoning benchmarks shows that RIFT consistently outperforms established baselines. This demonstrates that explicitly learning from mixed-quality data allows models to better internalize correct reasoning and failure modes. As a robust and data-efficient alignment method, RIFT enables scalable self-improvement without reliance on extensive expert-labeled data.
@@
Appendix A Extended Analysis and Empirical Explorations
In this section, we provide additional empirical evidence and in-depth analyses to further elucidate the behavior of RIFT. Specifically, we investigate how the scale of self-generated data influences the alignment performance and examine the underlying dynamics of policy convergence facilitated by our framework.
@@ A.1 Ablation Study on the Number of Self-Generated Responses
The performance trend is intriguing because the underlying data statistics, reported in Tables 12 and 13, show little variation across $K$: the proportion of positive responses remains stable at around 66.4%, and the fraction of problems with mixed reward signals plateaus near 85% for $K\geq 4$. In other words, simply generating more responses does not significantly alter the overall composition of the training data.
@@
The performance drop when increasing from $K=8$ to $K=16$ may stem from overfitting to noisy negative patterns in the limited MATH dataset. With more self-generated samples, the model encounters repeated or low-quality negatives, potentially leading to spurious correlations. This is consistent with observations in rejection sampling, where excessive sampling can degrade performance.
@@ A.2 Exploration Study: RIFT Drives Policy Convergence
As shown in Table 14, the results support the hypothesis that RIFT itself drives capability gains. Across all previously trained models, applying a second round of RIFT yields consistent improvements in Mean@3, outperforming their single-round counterparts: notable gains of +11.5 (SFT), +5.6 (DFT), and +1.2 (RIFT), even when starting from a strong RIFT-initialized policy. Gains diminish as the base policy strengthens, indicating convergence toward a shared high-reward policy. However, the re-application of RIFT primarily boosts Mean@3 rather than Pass@3, which shows little to no further improvement compared to single-round trained RIFT. Together, these findings demonstrate that RIFT acts not as a plug-and-play refiner whose efficacy depends on initialization, but as a self-convergent alignment phase that capable of steering diverse initial policies toward comparable final performance. Tables 15 and 16 provide a detailed statistical overview of the self-generated responses produced by the previously trained SFT, DFT, and RIFT models.
Appendix B Discussion on Loss Function and Gradient Stability
In this section, we provide a detailed analysis of our choice of using a Taylor approximation for the loss function, particularly in the context of penalizing negative samples.
B.1 Design Intuition: Stability over Global Precision
The primary motivation for employing a Taylor-based linear penalty (i.e., (1p)(1-p)) rather than the standard negative log-likelihood ($\log(1-p)$) for negative samples lies in the bounded and stable gradients it provides across the entire probability range $[0,1]$. While the Taylor approximation is locally accurate near $p\to 1$, its global behavior is empirically more conducive to stable optimization in model alignment.
For negative trajectories, our design intuition follows a specific priority:
  • Confident Errors: Tokens assigned high probability ($p\to 1$) on negative samples indicate cases where the model is “confidently wrong.” These require strong, linear penalties to steer the model away from incorrect reasoning paths.
  • Correct Uncertainty: Tokens with low probability ($p\to 0$) on negative examples reflect appropriate uncertainty. In such cases, the linear approximation provides a gentler correction signal, preventing the model from over-correcting on already unlikely tokens.
B.2 Comparison with Alternative Formulations
We evaluated several alternative functions to penalize negative samples, but found them less suitable for the following reasons:
  • Logarithmic Loss ($\log(1-p)$): This formulation diverges as $p\to 1$. In practical training, this leads to gradient instability and a significant magnitude imbalance between positive and negative loss components, often resulting in training divergence.
  • Sigmoid-based Variants: While sigmoid transformations can bound the loss, they introduce unnecessary saturation regions (where gradients vanish) without offering clear performance benefits in our preliminary experiments.
In summary, the choice of the Taylor approximation in RIFT is empirically motivated by its ability to maintain robust gradient signals and prevent optimization instability. We acknowledge that exploring other functional forms with similar bounded properties remains a promising direction for future work.
B.3 Proofs and Theoretical Analysis in Section 3
@@
Theorem B.1 (Formal Statement of Theorem 3.4).